OpenStudy (anonymous):

–3x – 8y = 24 ....How would i simplify this so that i get just y= mx +b so that i can graph this easier?

OpenStudy (kropot72):

Firstly add 3x to both sides of the equation to eliminate -3x from the left hand side. Can you do that?

OpenStudy (anonymous):

well yeah i get that part so -8 = 3x + 24

OpenStudy (anonymous):

-8y*

OpenStudy (kropot72):

Now divide each term on both sides of the equation by -8.

OpenStudy (anonymous):

ok so -y = ? +3 i dont know what i should do there? do I put it in decimal form?

OpenStudy (laddiusmaximus):

isnt this implicit differentation?

OpenStudy (kropot72):

\[\frac{-8y}{-8}=y\] \[\frac{3x+24}{-8}=-\frac{3}{8}x-3\]

OpenStudy (kropot72):

So the equation in the question converted to standard form becomes: \[y=-\frac{3}{8}x-3\]

OpenStudy (anonymous):

@TrogComplex you still graph it with y intercept ( 0, -3) and x intercept ( -8,0)